Quantity demanded= 20000-3P

quantity supplied= 15000+2P

a. calculate the equilibrium price and quantity.

b. calculate the price elasticity of supply using the point method when the economy is in equilibrium.



1
Expert's answer
2021-12-09T12:09:51-0500

a.

Qs=Qd

20000-3P=15000+2P

P=1000

Q=17000

b.

"Es=\frac{\frac{Q2-Q1}{\frac{Q2+Q1}{2}}}{\frac{P2-P1}{\frac{P2+P1}{2}}}"


P1=1000

P2=1200

Q1=17000

Q2=18000


"Es=\frac{\frac{Q2-Q1}{\frac{Q2+Q1}{2}}}{\frac{P2-P1}{\frac{P2+P1}{2}}}=\frac{\frac{18000-17000}{\frac{18000+17000}{2}}}{\frac{1200-1000}{\frac{1200+1000}{2}}}=\frac{0.057}{0.18}=0.31"